OpenStudy (anonymous):

Write an equation of a line parallel to line GH below in slope-intercept form that passes through the point (-5, 6).

OpenStudy (anonymous):

y = −x − 1 y = x − 1 y = −x + 11 y = x + 11

OpenStudy (anonymous):

@***[ISURU]***

OpenStudy (***[isuru]***):

can u find the intercept of graph ?